# Chain-linked Multiple Matrix Integration via Embedding Alignment

### v2

@@ 3 Theoretical Results
We now present theoretical guarantees for the estimate $\hat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}$ obtained by Algorithm 1. We shall make the following assumptions on the underlying population matrices $\{\mathbf{P}_{\mathcal{U}_{i},\mathcal{U}_{i}}\}$ for the observed blocks. We emphasize that, because our results address either large-sample approximations or limiting distributions, these assumptions should be interpreted in the regime where $n_{i}$ is arbitrarily large and/or $n_{i}\to\infty$.
@@
For the entire population matrix $\mathbf{P}\in\mathbb{R}^{N\times N}$, we have $\mathrm{rk}(\mathbf{P})=d$. Let $\lambda_{\max}$ and $\lambda_{\min}$ denote the largest and smallest non-zero eigenvalues of $\mathbf{P}$, and let $\mathbf{U}\in\mathbb{R}^{N\times d}$ contain the eigenvectors corresponding to non-zero eigenvalues. Suppose (1) $\mathbf{P}$ has bounded condition number, i.e., $\lambda_{\max}/\lambda_{\min}\leq M^{\prime}$ for some constant $M^{\prime}>0$, and $\mathbf{U}$ has bounded coherence, i.e., $\|\mathbf{U}\|_{2\to\infty}\lesssim d^{1/2}N^{-1/2}$; (2) for each $i$, $\mathcal{U}_{i}$ are drawn uniformly at random from $\mathcal{U}$. Then
@@
with high probability (Lemma D.11) and Eq. (3.1) holds.
We first consider the case where we only have two overlapping submatrices $\mathbf{A}^{(i)}$ and $\mathbf{A}^{(j)}$. Theorem 1 presents an expansion for $\hat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}-\mathbf{P}_{\mathcal{U}_{i},\mathcal{U}_{j}}=\hat{\mathbf{X}}^{(i)}\mathbf{W}^{(i,j)}\hat{\mathbf{X}}^{(j)\top}-\mathbf{X}_{\mathcal{U}_{i}}\mathbf{X}_{\mathcal{U}_{j}}^{\top}$.
@@
$n_{i,j}:=$ $|\mathcal{U}_{i}\cap\mathcal{U}_{j}|,\quad\vartheta_{i,j}:=\lambda_{\max}(\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}^{\top}\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}),\quad\theta_{i,j}:=\lambda_{\min}(\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}^{\top}\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}),$ (3.4)
$\alpha_{i,j}:=$ $\frac{n_{i,j}\gamma_{i}\gamma_{j}}{{\theta_{i,j}}(q_{i}n_{i}\mu_{i})^{1/2}(q_{j}n_{j}\mu_{j})^{1/2}}+\frac{{(n_{i,j}\vartheta_{i,j})^{1/2}}}{{\theta_{i,j}}}\Big{(}\frac{\gamma_{i}^{2}}{q_{i}n_{i}\mu_{i}^{3/2}}+\frac{\gamma_{j}^{2}}{q_{j}n_{j}\mu_{j}^{3/2}}\Big{)}$
@@
$\hat{\mathbf{P}}_{\mathcal{U}_{i},\mathcal{U}_{j}}-\mathbf{P}_{\mathcal{U}_{i},\mathcal{U}_{j}}$ $=\mathbf{E}^{(i)}\mathbf{X}_{\mathcal{U}_{i}}(\mathbf{X}_{\mathcal{U}_{i}}^{\top}\mathbf{X}_{\mathcal{U}_{i}})^{-1}\mathbf{X}_{\mathcal{U}_{j}}^{\top}+\mathbf{X}_{\mathcal{U}_{i}}(\mathbf{X}_{\mathcal{U}_{j}}^{\top}\mathbf{X}_{\mathcal{U}_{j}})^{-1}\mathbf{X}_{\mathcal{U}_{j}}^{\top}\mathbf{E}^{(j)}+\mathbf{R}^{(i,j)}+\mathbf{S}^{(i,j)},$ (3.5)
@@
$\alpha_{i,j}\min\Big{\{}\frac{(q_{i}n_{i}\mu_{i})^{1/2}}{\gamma_{i}}\|\mathbf{X}_{\mathcal{U}_{i}}\|_{2\to\infty},\frac{(q_{j}n_{j}\mu_{j})^{1/2}}{\gamma_{j}}\|\mathbf{X}_{\mathcal{U}_{j}}\|_{2\to\infty}\Big{\}}\lesssim 1.$
@@
$\|\hat{\mathbf{P}}_{\mathcal{U}_{i},\mathcal{U}_{j}}-\mathbf{P}_{\mathcal{U}_{i},\mathcal{U}_{j}}\|_{\max}\lesssim\frac{\gamma_{i}}{(q_{i}n_{i}\mu_{i})^{1/2}}\|\mathbf{X}_{\mathcal{U}_{j}}\|_{2\to\infty}+\frac{\gamma_{j}}{(q_{j}n_{j}\mu_{j})^{1/2}}\|\mathbf{X}_{\mathcal{U}_{i}}\|_{2\to\infty}$
@@
The expansion in Eq. (3.5) consists of four terms, with the first two terms being linear transformations of the additive noise matrices $\mathbf{E}^{(i)}$ and $\mathbf{E}^{(j)}$. The third term $\mathbf{R}^{(i,j)}$ corresponds to second-order estimation errors for $\hat{\mathbf{X}}^{(i)}$ and $\hat{\mathbf{X}}^{(j)}$, and hence Eq. (3.6) only depends on quantities associated with ${\mathbf{X}}_{\mathcal{U}_{i}}$ and ${\mathbf{X}}_{\mathcal{U}_{j}}$. Finally $\mathbf{S}^{(i,j)}$ corresponds to the error when aligning the overlaps $\hat{\mathbf{X}}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(i)}$ and $\hat{\mathbf{X}}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(j)}$ and hence Eq. (3.7) depends on $\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}$.
Remark 3.
Zhou et al. (2023) requires $(n_{i,j}/n_{i})\asymp(n_{i}/N)$ so that the number of overlapping entities must grow with the sizes of the submatrices. However, our derivations of Theorem 1 show that the overlap can be bounded or even be as small as $d$ (the rank of $\mathbf{P}$), which is the minimum value required to align the embeddings in $\mathbb{R}^{d}$; see Remark 6 and Section 4.3 for more detailed discussion and simulations verifying this result.
Next we consider a chain of overlapping submatrices as described in Algorithm 1. Theorem 2 presents the expansion of the estimation error for $\hat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}$ and Theorem 3 leverages this expansion to derive an entrywise normal approximation for $\hat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}-\mathbf{P}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}$. While the statements of Theorem 2 and Theorem 3 appear somewhat complicated at first glance, this is intentional as they make the results more general and thus applicable to a wider range of settings. Indeed, we allow for $(n_{i},\sigma_{i},q_{i})$ to have different magnitudes as well as the overlaps to be of very different sizes $n_{i,j}$. For example we can have $n_{1}\gg n_{2}\gg n_{3}$ but $q_{1}\ll q_{2}\ll q_{3}$ while $n_{12}\ll n_{23}$ but $\sigma_{1}\asymp\sigma_{2}\ll\sigma_{3}$. If $(n_{i},q_{i},\sigma_{i},n_{i,j})\equiv(n,q,\sigma,m)$ then these results can be simplified considerably; see Remark 6.
@@
$\hat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}-\mathbf{P}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}=\mathbf{E}^{({i_{0}})}\mathbf{X}_{\mathcal{U}_{i_{0}}}(\mathbf{X}_{\mathcal{U}_{i_{0}}}^{\top}\mathbf{X}_{\mathcal{U}_{i_{0}}})^{-1}\mathbf{X}_{\mathcal{U}_{i_{L}}}^{\top}+\mathbf{X}_{\mathcal{U}_{i_{0}}}(\mathbf{X}_{\mathcal{U}_{i_{L}}}^{\top}\mathbf{X}_{\mathcal{U}_{i_{L}}})^{-1}\mathbf{X}_{\mathcal{U}_{i_{L}}}^{\top}\mathbf{E}^{({i_{L}})}+\mathbf{R}^{(i_{0},i_{L})}+\mathbf{S}^{(i_{0},i_{1},\dots,i_{L})},$
@@
$\|\mathbf{R}^{(i_{0},i_{L})}\|_{\max}\lesssim$ $\Big{(}\frac{\gamma_{i_{0}}^{2}}{q_{i_{0}}n_{i_{0}}\mu_{i_{0}}^{3/2}}+\frac{\gamma_{i_{0}}}{q_{i_{0}}^{1/2}n_{i_{0}}\mu_{i_{0}}^{1/2}}\Big{)}\|\mathbf{X}_{\mathcal{U}_{i_{L}}}\|_{2\to\infty}+\Big{(}\frac{\gamma_{i_{L}}^{2}}{q_{i_{L}}n_{i_{L}}\mu_{i_{L}}^{3/2}}+\frac{\gamma_{i_{L}}}{q_{i_{L}}^{1/2}n_{i_{L}}\mu_{i_{L}}^{1/2}}\Big{)}\|\mathbf{X}_{\mathcal{U}_{i_{0}}}\|_{2\to\infty}$ (3.8)
$+\frac{\gamma_{i_{0}}\gamma_{i_{L}}}{(q_{i_{0}}n_{i_{0}}\mu_{i_{0}})^{1/2}(q_{i_{L}}n_{i_{L}}\mu_{i_{L}})^{1/2}},$
$\|\mathbf{S}^{(i_{0},i_{1},\dots,i_{L})}$ $\|_{\max}\lesssim\Big{[}\sum_{\ell=1}^{L}{\alpha_{{i_{\ell-1}},{i_{\ell}}}}\Big{]}\cdot\|\mathbf{X}_{\mathcal{U}_{i_{0}}}\|_{2\to\infty}\cdot\|\mathbf{X}_{\mathcal{U}_{i_{L}}}\|_{2\to\infty}$
@@
$\Big{[}\sum_{\ell=1}^{L}{\alpha_{{i_{\ell-1}},{i_{\ell}}}}\Big{]}\min\Big{\{}\frac{(q_{i_{0}}n_{i_{0}}\mu_{i_{0}})^{1/2}}{\gamma_{i_{0}}}\|\mathbf{X}_{\mathcal{U}_{i_{0}}}\|_{2\to\infty},\frac{(q_{i_{L}}n_{i_{L}}\mu_{i_{L}})^{1/2}}{\gamma_{i_{L}}}\|\mathbf{X}_{\mathcal{U}_{i_{L}}}\|_{2\to\infty}\Big{\}}\lesssim 1.$
@@
$\|\hat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}-\mathbf{P}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}\|_{\max}\lesssim\frac{\gamma_{i_{0}}}{(q_{i_{0}}n_{i_{0}}\mu_{i_{0}})^{1/2}}\|\mathbf{X}_{\mathcal{U}_{i_{L}}}\|_{2\to\infty}+\frac{\gamma_{i_{L}}}{(q_{i_{L}}n_{i_{L}}\mu_{i_{L}})^{1/2}}\|\mathbf{X}_{\mathcal{U}_{i_{0}}}\|_{2\to\infty}$
@@
We note that the only difference between Theorem 1 and Theorem 2 is in the upper bound for $\mathbf{S}^{(i_{0},i_{1},\dots,i_{L})}$ compared to that for $\mathbf{S}^{(i,j)}$. Indeed, $\mathbf{E}^{({i_{0}})}\mathbf{X}_{\mathcal{U}_{i_{0}}}(\mathbf{X}_{\mathcal{U}_{i_{0}}}^{\top}\mathbf{X}_{\mathcal{U}_{i_{0}}})^{-1}\mathbf{X}_{\mathcal{U}_{i_{L}}}^{\top}+\mathbf{X}_{\mathcal{U}_{i_{0}}}(\mathbf{X}_{\mathcal{U}_{i_{L}}}^{\top}\mathbf{X}_{\mathcal{U}_{i_{L}}})^{-1}\mathbf{X}_{\mathcal{U}_{i_{L}}}^{\top}\mathbf{E}^{({i_{L}})}$ and $\mathbf{R}^{(i_{0},i_{L})}$ in Theorem 2 only depend on $\mathbf{X}_{\mathcal{U}_{i_{0}}}$ and $\mathbf{X}_{\mathcal{U}_{i_{L}}}$, but not on the chain linking them, and thus their upper bounds are the same as that in Theorem 1 for $i=i_{0}$ and $j=i_{L}$. In contrast, from our discussion in Remark 2, the term $\mathbf{S}^{(i_{0},i_{1},\dots,i_{L})}$ corresponds to the alignment error between $\hat{\mathbf{X}}^{(i_{0})}$ and $\hat{\mathbf{X}}^{(i_{L})}$. As $\mathcal{U}_{i_{0}}$ and $\mathcal{U}_{i_{L}}$ need not share any overlap, this alignment is obtained via a sequence of orthogonal Procrustes transformations between $\hat{\mathbf{X}}^{(i_{\ell-1})}_{\langle\mathcal{U}_{i_{\ell-1}}\cap\mathcal{U}_{i_{\ell}}\rangle}$ and $\hat{\mathbf{X}}^{(i_{\ell})}_{\langle\mathcal{U}_{i_{\ell-1}}\cap\mathcal{U}_{i_{\ell}}\rangle}$ for $1\leq\ell\leq L$. The accumulated error for these transformations is reflected in the term $\sum_{\ell=1}^{L}{\alpha_{i_{\ell-1},i_{\ell}}}$ and depends on the whole chain. When the length of the chain $L$ is not too large relative to the overlaps, the error in $\mathbf{S}^{(i_{0},i_{1},\dots,i_{L})}$ is negligible compared to that in the dominant term $\mathbf{E}^{({i_{0}})}\mathbf{X}_{\mathcal{U}_{i_{0}}}(\mathbf{X}_{\mathcal{U}_{i_{0}}}^{\top}\mathbf{X}_{\mathcal{U}_{i_{0}}})^{-1}\mathbf{X}_{\mathcal{U}_{i_{L}}}^{\top}+\mathbf{X}_{\mathcal{U}_{i_{0}}}(\mathbf{X}_{\mathcal{U}_{i_{L}}}^{\top}\mathbf{X}_{\mathcal{U}_{i_{L}}})^{-1}\mathbf{X}_{\mathcal{U}_{i_{L}}}^{\top}\mathbf{E}^{({i_{L}})}$, and consequently, our entrywise error rate depends only on $\mathbf{X}_{\mathcal{U}_{i_{0}}}$ and $\mathbf{X}_{\mathcal{U}_{i_{L}}}$, rather than on the chain linking them.

### v4

@@ 3 Theoretical Results
We now present theoretical guarantees for the estimate $\widehat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}$ obtained by Algorithm 1. We shall make the following assumptions on the underlying population matrices $\{\mathbf{P}_{\mathcal{U}_{i},\mathcal{U}_{i}}\}$ for the observed blocks. We emphasize that, because our results address either large-sample approximations or limiting distributions, these assumptions should be interpreted in the regime where $n_{i}$ is arbitrarily large and/or $n_{i}\to\infty$.
@@
We note that the conditions in Assumption 1 are quite mild and typically seen in the matrix completion literature. For example Eq. (3.1) is satisfied whenever $\mathbf{P}_{\mathcal{U}_{i},\mathcal{U}_{i}}$ has bounded condition number and bounded coherence; see e.g., Abbe et al. (2020); Chen et al. (2021); Recht (2011). Next, $n_{i}q_{i}\gtrsim\log^{2}{n_{i}}$ is much less stringent compared to $q_{i}\equiv 1$ as assumed in Zhou et al. (2023) and Bishop and Yu (2014). Finally Eq. (3.2) is satisfied whenever Eq. (3.1) holds and $\|\mathbf{P}_{\mathcal{U}_{i},\mathcal{U}_{i}}\|_{\max}+\sigma_{i}=O(1)$. See Remark 5 for further discussion. Note that $\|\mathbf{P}_{\mathcal{U}_{i},\mathcal{U}_{i}}\|_{\max}=\|\mathbf{X}_{\mathcal{U}_{i}}\|_{2\to\infty}^{2}\lesssim d\frac{\lambda_{i,\max}}{n_{i}}\asymp\mu_{i}$ under Assumption 1, and see Section E.9 of the supplementary material for further discussion on $\mu_{i}$ and $\|\mathbf{X}_{\mathcal{U}_{i}}\|_{2\to\infty}$.
@@
Let $\mathbf{P}\in\mathbb{R}^{N\times N}$ have rank $d$. Denote by $\lambda_{\max}$ and $\lambda_{\min}$ the largest and smallest non-zero eigenvalues of $\mathbf{P}$, respectively, and let $\mathbf{U}\in\mathbb{R}^{N\times d}$ contain the eigenvectors corresponding to the non-zero eigenvalues of $\mathbf{P}$. Suppose (1) $\mathbf{P}$ has bounded condition number, i.e., $\lambda_{\max}/\lambda_{\min}\leq M^{\prime}$ for some constant $M^{\prime}>0$, and $\mathbf{U}$ has bounded coherence, i.e., $\|\mathbf{U}\|_{2\to\infty}\lesssim d^{1/2}N^{-1/2}$; (2) for each $i$, $\mathcal{U}_{i}$ are drawn uniformly at random from $\mathcal{U}$. Then
@@
with high probability, and Eq. (3.1) holds (see Lemma E.11 for more details).
We first present an informal, simplified version of the theoretical results to provide a basic intuition about the properties of CMMI, as the formal statement, while being quite more general, is also more complex. Suppose that $n_{i}\asymp n,q_{i}\equiv q,\sigma_{i}\equiv\sigma$, and $\lambda_{i,\min}\asymp\frac{n_{i}}{N}\lambda_{\min}$ for all $i$, with the overlap sizes satisfying $n_{i_{\ell-1},i_{\ell}}\equiv m\geq d$ for all $1\leq\ell\leq L$. This setting corresponds to a scenario where the block sizes, overlap sizes, noise levels, and missing rates are uniform, and we further suppose that the entries are bounded (see Remark 5 for more details on this case). Then we have the entrywise error bound
$\|\widehat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}-\mathbf{P}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}\|_{\max}\lesssim\frac{(1+\sigma)\log^{1/2}n}{(qn)^{1/2}}$
with high probability, provided that $L\Big{(}\frac{(1+\sigma)\log^{1/2}n}{(qn)^{1/2}}+\frac{1}{m^{1/2}}\Big{)}\lesssim 1$. Furthermore, under mild conditions we also establish an entrywise normal approximation. More specifically, for any $s\in[n_{i_{0}}],t\in[n_{i_{L}}]$, we have
$\widetilde{\sigma}_{s,t}^{-1}\big{(}\widehat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}-\mathbf{P}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}\big{)}_{s,t}\rightsquigarrow\mathcal{N}(0,1)$
as $n\to\infty$, where the standard deviation satisfies $\widetilde{\sigma}_{s,t}\lesssim\frac{\sigma^{2}+(1-q)}{qn}.$
We now present the formal theoretical results, which are applicable to a much broader range of settings while also yielding more detailed analysis. We first consider the case where we only have two overlapping submatrices $\mathbf{A}^{(i)}$ and $\mathbf{A}^{(j)}$. Theorem 1 presents an expansion for $\widehat{\mathbf{P}}_{\mathcal{U}_{i},\mathcal{U}_{j}}-\mathbf{P}_{\mathcal{U}_{i},\mathcal{U}_{j}}=\widehat{\mathbf{X}}^{(i)}\mathbf{W}^{(i,j)}\widehat{\mathbf{X}}^{(j)\top}-\mathbf{X}_{\mathcal{U}_{i}}\mathbf{X}_{\mathcal{U}_{j}}^{\top}$.
@@
$n_{i,j}$= $|\mathcal{U}_{i}\cap\mathcal{U}_{j}|,\quad\vartheta_{i,j}=\lambda_{\max}(\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}^{\top}\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}),\quad\theta_{i,j}=\lambda_{\min}(\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}^{\top}\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}),$ (3.4)
$\alpha_{i,j}$= $\frac{n_{i,j}\gamma_{i}\gamma_{j}}{{\theta_{i,j}}(q_{i}n_{i}\mu_{i})^{1/2}(q_{j}n_{j}\mu_{j})^{1/2}}+\frac{{(n_{i,j}\vartheta_{i,j})^{1/2}}}{{\theta_{i,j}}}\Big{(}\frac{\gamma_{i}^{2}}{q_{i}n_{i}\mu_{i}^{3/2}}+\frac{\gamma_{j}^{2}}{q_{j}n_{j}\mu_{j}^{3/2}}\Big{)}$
@@
$\widehat{\mathbf{P}}_{\mathcal{U}_{i},\mathcal{U}_{j}}-\mathbf{P}_{\mathcal{U}_{i},\mathcal{U}_{j}}$ $=\mathbf{E}^{(i)}\mathbf{X}_{\mathcal{U}_{i}}(\mathbf{X}_{\mathcal{U}_{i}}^{\top}\mathbf{X}_{\mathcal{U}_{i}})^{-1}\mathbf{X}_{\mathcal{U}_{j}}^{\top}+\mathbf{X}_{\mathcal{U}_{i}}(\mathbf{X}_{\mathcal{U}_{j}}^{\top}\mathbf{X}_{\mathcal{U}_{j}})^{-1}\mathbf{X}_{\mathcal{U}_{j}}^{\top}\mathbf{E}^{(j)}+\mathbf{R}^{(i,j)}+\mathbf{S}^{(i,j)},$ (3.5)
@@
$\alpha_{i,j}\min\Big{\{}\frac{(q_{i}n_{i}\mu_{i})^{1/2}}{\gamma_{i}}\|\mathbf{X}_{\mathcal{U}_{i}}\|_{2\to\infty},\frac{(q_{j}n_{j}\mu_{j})^{1/2}}{\gamma_{j}}\|\mathbf{X}_{\mathcal{U}_{j}}\|_{2\to\infty}\Big{\}}\lesssim 1.$ (3.8)
@@
$\|\widehat{\mathbf{P}}_{\mathcal{U}_{i},\mathcal{U}_{j}}-\mathbf{P}_{\mathcal{U}_{i},\mathcal{U}_{j}}\|_{\max}\lesssim\frac{\gamma_{i}}{(q_{i}n_{i}\mu_{i})^{1/2}}\|\mathbf{X}_{\mathcal{U}_{j}}\|_{2\to\infty}+\frac{\gamma_{j}}{(q_{j}n_{j}\mu_{j})^{1/2}}\|\mathbf{X}_{\mathcal{U}_{i}}\|_{2\to\infty}$
@@
The expansion in Eq. (3.5) consists of four terms, with the first two terms being linear transformations of the additive noise matrices $\mathbf{E}^{(i)}$ and $\mathbf{E}^{(j)}$. The third term $\mathbf{R}^{(i,j)}$ corresponds to second-order estimation errors for $\widehat{\mathbf{X}}^{(i)}$ and $\widehat{\mathbf{X}}^{(j)}$, and hence Eq. (3.6) only depends on quantities associated with ${\mathbf{X}}_{\mathcal{U}_{i}}$ and ${\mathbf{X}}_{\mathcal{U}_{j}}$. The last term $\mathbf{S}^{(i,j)}$ corresponds to the error when aligning the overlaps $\widehat{\mathbf{X}}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(i)}$ and $\widehat{\mathbf{X}}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}^{(j)}$ and hence Eq. (3.7) depends on $\mathbf{X}_{\mathcal{U}_{i}\cap\mathcal{U}_{j}}$. Finally, Eq. (3.8) ensures that $\mathbf{S}^{(i,j)}$ is bounded by the first two terms, and is a mild and natural assumption in many settings; see Remark 5 for further discussion.
Next we consider a chain of overlapping submatrices as described in Algorithm 1. Theorem 2 presents the expansion of the estimation error for $\widehat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}$ and Theorem 3 leverages this expansion to derive an entrywise normal approximation for $\widehat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}-\mathbf{P}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}$. While the statements of Theorem 2 and Theorem 3 appear somewhat complicated at first glance, this is intentional as they make the results more general and thus applicable to a wider range of settings. Indeed, we allow for $(n_{i},\sigma_{i},q_{i})$ to have different magnitudes as well as the overlaps to be of very different sizes $n_{i,j}$. For example we can have $n_{1}\gg n_{2}\gg n_{3}$ but $q_{1}\ll q_{2}\ll q_{3}$ while $n_{1,2}\ll n_{2,3}$ but $\sigma_{1}\asymp\sigma_{2}\ll\sigma_{3}$. If $(n_{i},q_{i},\sigma_{i},n_{i,j})\equiv(n,q,\sigma,m)$ then these results can be simplified considerably; see Remark 5.
@@
$\widehat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}-\mathbf{P}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}=\mathbf{E}^{({i_{0}})}\mathbf{X}_{\mathcal{U}_{i_{0}}}(\mathbf{X}_{\mathcal{U}_{i_{0}}}^{\top}\mathbf{X}_{\mathcal{U}_{i_{0}}})^{-1}\mathbf{X}_{\mathcal{U}_{i_{L}}}^{\top}+\mathbf{X}_{\mathcal{U}_{i_{0}}}(\mathbf{X}_{\mathcal{U}_{i_{L}}}^{\top}\mathbf{X}_{\mathcal{U}_{i_{L}}})^{-1}\mathbf{X}_{\mathcal{U}_{i_{L}}}^{\top}\mathbf{E}^{({i_{L}})}+\mathbf{R}^{(i_{0},i_{L})}+\mathbf{S}^{(i_{0},i_{1},\dots,i_{L})},$
@@
$\begin{split}\|\mathbf{R}^{(i_{0},i_{L})}\|_{\max}\lesssim&\frac{\gamma_{i_{0}}\gamma_{i_{L}}}{(q_{i_{0}}n_{i_{0}}\mu_{i_{0}})^{1/2}(q_{i_{L}}n_{i_{L}}\mu_{i_{L}})^{1/2}}+\Big{(}\frac{\gamma_{i_{0}}^{2}}{q_{i_{0}}n_{i_{0}}\mu_{i_{0}}^{3/2}}+\frac{\gamma_{i_{0}}}{q_{i_{0}}^{1/2}n_{i_{0}}\mu_{i_{0}}^{1/2}}\Big{)}\|\mathbf{X}_{\mathcal{U}_{i_{L}}}\|_{2\to\infty}\\ &+\Big{(}\frac{\gamma_{i_{L}}^{2}}{q_{i_{L}}n_{i_{L}}\mu_{i_{L}}^{3/2}}+\frac{\gamma_{i_{L}}}{q_{i_{L}}^{1/2}n_{i_{L}}\mu_{i_{L}}^{1/2}}\Big{)}\|\mathbf{X}_{\mathcal{U}_{i_{0}}}\|_{2\to\infty}\end{split}$ (3.9)
$\|\mathbf{S}^{(i_{0},i_{1},\dots,i_{L})}\|_{\max}\lesssim\Big{[}\sum_{\ell=1}^{L}{\alpha_{{i_{\ell-1}},{i_{\ell}}}}\Big{]}\cdot\|\mathbf{X}_{\mathcal{U}_{i_{0}}}\|_{2\to\infty}\cdot\|\mathbf{X}_{\mathcal{U}_{i_{L}}}\|_{2\to\infty}$ (3.10)
@@
$\Big{[}\sum_{\ell=1}^{L}{\alpha_{{i_{\ell-1}},{i_{\ell}}}}\Big{]}\min\Big{\{}\frac{(q_{i_{0}}n_{i_{0}}\mu_{i_{0}})^{1/2}}{\gamma_{i_{0}}}\|\mathbf{X}_{\mathcal{U}_{i_{0}}}\|_{2\to\infty},\frac{(q_{i_{L}}n_{i_{L}}\mu_{i_{L}})^{1/2}}{\gamma_{i_{L}}}\|\mathbf{X}_{\mathcal{U}_{i_{L}}}\|_{2\to\infty}\Big{\}}\lesssim 1.$ (3.11)
@@
$\|\widehat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}-\mathbf{P}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}\|_{\max}\lesssim\frac{\gamma_{i_{0}}}{(q_{i_{0}}n_{i_{0}}\mu_{i_{0}})^{1/2}}\|\mathbf{X}_{\mathcal{U}_{i_{L}}}\|_{2\to\infty}+\frac{\gamma_{i_{L}}}{(q_{i_{L}}n_{i_{L}}\mu_{i_{L}})^{1/2}}\|\mathbf{X}_{\mathcal{U}_{i_{0}}}\|_{2\to\infty}$
@@
The only difference between Theorems 1 and 2 is in the upper bound for $\mathbf{S}^{(i_{0},i_{1},\dots,i_{L})}$ compared to that for $\mathbf{S}^{(i,j)}$. Indeed, $\mathbf{E}^{({i_{0}})}\mathbf{X}_{\mathcal{U}_{i_{0}}}(\mathbf{X}_{\mathcal{U}_{i_{0}}}^{\top}\mathbf{X}_{\mathcal{U}_{i_{0}}})^{-1}\mathbf{X}_{\mathcal{U}_{i_{L}}}^{\top}+\mathbf{X}_{\mathcal{U}_{i_{0}}}(\mathbf{X}_{\mathcal{U}_{i_{L}}}^{\top}\mathbf{X}_{\mathcal{U}_{i_{L}}})^{-1}\mathbf{X}_{\mathcal{U}_{i_{L}}}^{\top}\mathbf{E}^{({i_{L}})}$ and $\mathbf{R}^{(i_{0},i_{L})}$ in Theorem 2 only depend on $\mathbf{X}_{\mathcal{U}_{i_{0}}}$ and $\mathbf{X}_{\mathcal{U}_{i_{L}}}$, but not on the chain linking them, and thus their upper bounds are the same as that in Theorem 1 for $i=i_{0}$ and $j=i_{L}$. In contrast, from our discussion in Remark 2, the term $\mathbf{S}^{(i_{0},i_{1},\dots,i_{L})}$ corresponds to the alignment error between $\widehat{\mathbf{X}}^{(i_{0})}$ and $\widehat{\mathbf{X}}^{(i_{L})}$. As $\mathcal{U}_{i_{0}}$ and $\mathcal{U}_{i_{L}}$ need not share any overlap, this alignment is obtained via a sequence of orthogonal Procrustes transformations between $\widehat{\mathbf{X}}^{(i_{\ell-1})}_{\langle\mathcal{U}_{i_{\ell-1}}\cap\mathcal{U}_{i_{\ell}}\rangle}$ and $\widehat{\mathbf{X}}^{(i_{\ell})}_{\langle\mathcal{U}_{i_{\ell-1}}\cap\mathcal{U}_{i_{\ell}}\rangle}$ for $1\leq\ell\leq L$. The accumulated error for these transformations is reflected in the term $\sum_{\ell=1}^{L}{\alpha_{i_{\ell-1},i_{\ell}}}$ and depends on the whole chain. If $L$ is not too large relative to the overlaps $\{n_{i,j}\}$ then the error in $\mathbf{S}^{(i_{0},i_{1},\dots,i_{L})}$ is negligible compared to that of $\mathbf{E}^{({i_{0}})}\mathbf{X}_{\mathcal{U}_{i_{0}}}(\mathbf{X}_{\mathcal{U}_{i_{0}}}^{\top}\mathbf{X}_{\mathcal{U}_{i_{0}}})^{-1}\mathbf{X}_{\mathcal{U}_{i_{L}}}^{\top}+\mathbf{X}_{\mathcal{U}_{i_{0}}}(\mathbf{X}_{\mathcal{U}_{i_{L}}}^{\top}\mathbf{X}_{\mathcal{U}_{i_{L}}})^{-1}\mathbf{X}_{\mathcal{U}_{i_{L}}}^{\top}\mathbf{E}^{({i_{L}})}$, and consequently our entrywise error rate depends only on $\mathbf{X}_{\mathcal{U}_{i_{0}}}$ and $\mathbf{X}_{\mathcal{U}_{i_{L}}}$ rather than on the chain linking them.